OpenStudy (anonymous):

convert these to radians please please -230 degrees -280 degrees -30 degrees 514 degrees

OpenStudy (anonymous):

once again you have to use the formula degrees(pi/180) to get radian -230=-23pi/18 -280=-14/9 -30=-pi/6 514= 257pi/90

OpenStudy (anonymous):

514 does not work

OpenStudy (anonymous):

\[514 \times \pi / 180 \]=257pi/90